Two thirds of what number is equal to 18 plus 60% of 40?
(A) 27
(B) 33
(C) 36
(D) 63

Answer by rfer(16322)

2/3x=40*0.6+18
2/3x=42
x=42/(2/3)
x=42*3/2
x=63